OpenStudy (anonymous):

Can anyone help with algebra 1

OpenStudy (anonymous):

Heres the question Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 2^-x and y = 4^x +^ 3 intersect are the solutions of the equation 2^-x = 4^x + ^3. Part B: Make tables to find the solution to 2^-x = 4^x +^ 3. Take the integer values of x only between -3 and 3. Part C: How can you solve the equation 2^-x = 4^x + ^3 graphically?

OpenStudy (anonymous):

i will medal